# Cascade switching current detectors based on arrays of Josephson junctions

### v1

@@ Discussion
As described in the Supplementary, for Bi-2212 detector, the signal, $\Delta V$, increases from 0.32 mV for the single junction SCD (no amplification, $n=1$), to 81 mV at sub-optimal bias, and almost up to 0.3 V at the optimal bias, Fig. 2 (e). We emphasize that these measurements were performed under the same conditions with the only difference in $I_{b}$, determining the cascade amplification.
@@
The sensitivity $S$(V/W) can be calculated in the following way. The readout voltage from the array is $V(n)=nV_{c}$. The spread of switching currents for JJs in the array can be written as, $I_{s}(n)\simeq I_{s}^{*}+\sigma_{n}(n/N-1/2)$, where $I_{s}^{*}$ is the mean switching current and $\sigma_{n}$ is the width of the multi-junction switching current distribution. At the optimal bias, $I_{b}\simeq I_{s}^{*}$, the SCD response, Eq. (1), is $\Delta V\sim NV_{c}(I_{THz}/\sigma_{n})$. The absorbed THz power can be writen as, $P_{THz}=I_{THz}^{2}R_{THz}/2$, where $R_{THz}$ is the real part of the high-frequency impedance of the detector. Thus, we obtain
$S\simeq\frac{2NV_{c}}{\sigma_{n}R_{THz}I_{THz}}.$ (2)
@@
The procedure for quantitative estimation of sensitivity and noise-equivalent power (NEP) is described in Methods. For the case of Bi-2212 detectors from Figs. 2 and 3 we get $S>10^{7}$ (V/W) and $NEP\sim 10^{-17}~{}$(W$/\sqrt{Hz}$) for both MW and THz detection. There are very good numbers, however, far from optimal. Lets estimate achievable characteristics for a high-quality Bi-2212 mesa, like in Fig. 2 (c). Assuming $N=500$, $V_{c}=20$ mV, $I_{s}^{*}=50~{}\upmu$A; $\sigma_{n}=0.01I_{s}^{*}$ (typical for fluctuation-induced $\sigma_{n}$ at $T\sim 4$ K [55, 13]); $R_{THz}=100~{}\Omega$ (for an efficient detector $R_{THz}$ should be comparable to the impedance of free space); and $I_{THz}=0.1\sigma_{n}=50~{}$nA; we obtain $S\simeq 8\cdot 10^{12}$ (V/W) and $NEP\simeq 10^{-22}~{}$(W$/\sqrt{Hz}$). This estimation is made under realistic assumptions and does not represent the ultimate limit of SCD [13].
@@
Two factors contribute to this high sensitivity: the large cascade gain $\propto N$ and the narrow switching width $\sigma_{n}$. The latter depends on the uniformity of the array, inter-junction coupling, thermal fluctuations and electrical noise. When all junction in the stack are identical, they will switch at the same bias current. A spread of parameters cause broadening of $\sigma_{n}$. However, coupling between JJs leads to current-locking even of slightly dissimilar junctions [41]. The true phase-locking of JJs can further reduce $\sigma_{n}$ [57]. Therefore, a strong inter-junction coupling is crucial for such a detector. Ultimately, $\sigma_{n}$ is limited by quantum fluctuations [19, 21, 22, 23, 29, 55], qualifying SCD as a quantum-limited photon detector.
@@
A good detector should possess both high sensitivity to absorbed power and high absorption efficiency of incident radiation. The maximum efficiency ($50\%$) is achieved at the impedance matching condition [46, 45, 44]. However, a single JJ alone cannot provide proper matching because it is much smaller than $\lambda_{0}$ [44]. Therefore, the geometry of electrodes forming a receiving antenna becomes crucial. In the case of linear Nb arrays, long electrodes function as traveling wave antennas, as evidenced by off-axis diagrams of both emission [50, 51] and absorption, Fig. 1 (f). Bi-2212 whisker mesas have a turnstile antenna geometry, enabling good impedance matching [40]. This is evidenced by the high radiation efficiency ($12\%$ at $f\simeq 4$ THz) [27] and the four-fold polarization diagram reported in Fig. 2 (f) [56]. Thus, the analysis of geometry-specific antenna effects provides a valuable tool for investigating impedance matching and the efficiency of both absorption and emission.
@@
In conclusion, we have described the operational principle of a cascade switching current detector, utilizing arrays of coupled Josephson junctions. Coupling between junctions leads to the current locking phenomenon and results in avalanche-like switching of many junctions. This amplifies the readout voltage without changing the noise background and, therefore, enhances the signal-to-noise ratio and reduces NEP. We have demonstrated two prototypes based either on low-$T_{c}$ linear Nb arrays, or high-$T_{c}$ stacked intrinsic Josephson junctions. Good impedance matching and absorption efficiency of the studied devices is confirmed by observation of geometry-specific antenna effects. We conclude that Bi-2212 whisker-based cascade SCD are promising candidates for broad-band and high-performance THz sensors.
@@ Methods
Experimental details
Measurements are performed in a closed-cycle optical cryostat with a base temperature $\sim 3$ K at zero magnetic field. Arrays are current-biased using a programmable voltage source. The $I$-$V$ characteristics are measured in a quasi 4-probe configuration. Additional information about experimental setup, bias configuration and microwave measurements can be found in sec. SVII of the Supplementary. Details about fabrication and characterization of Nb arrays and Bi-2212 mesas are provided in sec. SVIII and SIX.
Operation of SCD requires statistical analysis. The statistic of Bi-2212 mesas for a single and multi-junction switching can be found in sec. SIV of the Supplementary. It is well described by the conventional thermal-activation theory, described in sec. SII.
The SCD response is measured via lock-in measurements. For this, a harmonic current, $I=I_{b}\sin(2\pi f_{b}t)$, with the amplitude $I_{b}$ and frequency $f_{b}=23$ Hz is supplied for a time interval of 1 s. The read-out voltage corresponds to the Fourier component of the detector $V(t)$ wave form at $f=f_{b}$. The lock-in response is calculated on-flight using FPGA-based FFT routine. Additional information about SCD operation and lock-in readout can be found in sec. SI and SV of the Supplementary.
@@
The absorbed power, $P_{a}$, can be estimated from the suppression of the switching current, $\Delta I_{s}$, caused by the generated high-frequency current. In the MW case (low frequency limit) $\Delta I_{s}$ is just equal to $I_{MW}$ [13]. $P_{a}$ is calculated as [24]: $P_{a}\simeq(2\sqrt{2}/3\pi)(\Delta I_{s}/I_{c0})^{3/2}I_{c0}V_{c}$. For the case of Fig. 2 (d), $\Delta I_{s}\simeq 3~{}\mu$A at the attenuation 3.7 dB, $I_{c0}=38~{}\mu$A, $V_{c}=20~{}$mV, yielding $P_{a}\simeq 5$ nW, approximately 30% of $P_{MW}$. The maximum voltage is $V\simeq 0.37$ V, and the sensitivity with respect to the absorbed power, $S_{a}=V/P_{a}\simeq 7.4~{}10^{7}$ (V/W) and the net sensitivity $S_{MW}=V/P_{MW}\simeq 2.6~{}10^{7}$ (V/W).
@@
The noise-equivalent power is determined as $NEP=v_{n}/S$, where $v_{n}$ is the voltage noise. As shown in the Supplementary, in our setup it is in the range of few nV$/\sqrt{Hz}$, but could be reduced to below 1 nV$/\sqrt{Hz}$ using cryogenic preamplifiers. Taking $v_{n}=1~{}$nV$/\sqrt{Hz}$ we obtain the following values of NEP with respect to absorbed power: $NEP_{MW}\simeq 1.35~{}10^{-17}$ W$/\sqrt{Hz}$ and $NEP_{THz}\simeq 6.25~{}10^{-17}$ W$/\sqrt{Hz}$.
@@
Noise characteristics
The signal-to-noise ratio in cascade SCD is determined by the three quantities:
- The mean readout voltage $V_{n}$,
- The standard deviation (statistical uncertainty) $\delta V_{n}$,
- The measurement noise threshold.
Direct noise measurements in our setup are presented in sec. SVI of the Supplementary. The noise floor is in the range 1-10 nV/Hz1/2, limited by the intrinsic noise in the room temperature preamplifier.
The effect of statistical uncertainty and the difference in sensitivities of single junction SCD and cascade SCD are discussed in sec. SIII. There we also write explicit equations for the sensitivity and provide corresponding numerical simulations.
The bottom line of this analysis is the following: Cascading increases the mean readout $V_{n}$, generally improves the relative uncertainty, $\delta V_{n}/V_{n}$, and usually expands the dynamic range. But it does not affect the measurement noise (determined by the preamplifier). Therefore, cascading always increases the signal-to-noise ratio and reduces NEP, compared to a single-junction SCD.
@@ Acknowledgements
The work was supported by the Science for Peace and Security Programme, grant G5796. We are grateful to R. Gerdau and M. Galin for assistance with the fabrication of Nb-arrays and to A. Kalenyuk for assistance with the fabrication of Bi-2212 samples.

### v2

@@ Discussion
[…]
@@
The detector sensitivity $S$(V/W) can be calculated in the following way. The readout voltage from the array is $V(n)=nV_{c}$. The spread of switching currents for JJs in the array can be written as, $I_{s}(n)\simeq I_{s}^{*}+\sigma(n/N-1/2)$, where $I_{s}^{*}$ is the mean switching current and $\sigma$ is the width of the switching current distribution. At the optimal bias, $I_{b}\simeq I_{s}^{*}$, the effective differential resistance is, $R_{d}\simeq NV_{c}/\sigma$. The response of SCD is $\Delta V\simeq R_{d}I_{THz}$, where $I_{THz}$ is the induced THz current. The absorbed THz power can be writen as, $P_{THz}=I_{THz}^{2}R_{THz}/2$, where $R_{THz}$ is the real part of the high-frequency impedance of the detector (including electrodes). Thus, we obtain
$S\simeq\frac{2NV_{c}}{\sigma R_{THz}I_{THz}}.$ (1)
@@
The procedure for quantitative estimation of sensitivity and noise-equivalent power (NEP) is described in Methods. For the case of Bi-2212 detectors from Figs. 2 and 3 we get $S>10^{7}$ (V/W) and $NEP\sim 10^{-17}~{}$(W$/\sqrt{Hz}$) for both MW and THz detection. There are very good numbers, however, far from optimal. Lets estimate achievable characteristics for a small, high-quality Bi-2212 mesa, like in Fig. 2 (c). Assuming $N=500$, $V_{c}=20$, $I_{s}^{*}=50~{}\upmu$A; $\sigma=0.01I_{s}^{*}$ (typical for fluctuation-induced $\sigma$ at $T\sim 4$ K [49, 8]); $R_{THz}=100~{}\Omega$ (for an efficient detector $R_{THz}$ should be comparable to the impedance of free space); and $I_{THz}=0.1\sigma=50~{}$nA; we obtain $S\simeq 8\cdot 10^{12}$ (V/W) and $NEP\simeq 10^{-22}~{}$(W$/\sqrt{Hz}$). This estimation is made under realistic assumptions and does not represent the ultimate limit of SCD [8].
@@
Two factors contribute to this high sensitivity: the large cascade gain $\propto N$ and the narrow switching width $\sigma$. The latter depends on the uniformity of the array, inter-junction coupling, thermal fluctuations and electrical noise. When all junction in the stack are identical, they will switch at the same bias current. A spread of parameters cause broadening of $\sigma$. However, coupling between JJs leads to current-locking even of slightly dissimilar junctions [36]. The true phase-locking of JJs can further reduce $\sigma$ [51]. Therefore, a strong inter-junction coupling is crucial for such a detector. Ultimately, $\sigma$ is limited by quantum fluctuations [14, 16, 17, 18, 24, 49], qualifying SCD as a quantum-limited photon detector.
@@
A good detector should possess both high sensitivity to absorbed power and high absorption efficiency of incident radiation. The maximum efficiency ($50\%$) is achieved at the impedance matching condition [41, 40, 39]. However, a single JJ alone cannot provide proper matching because it is much smaller than $\lambda_{0}$ [39]. Therefore, the geometry of electrodes forming a receiving antenna becomes crucial. In the case of linear Nb arrays, long electrodes function as traveling wave antennas, as evidenced by off-axis directivities of both emission [45, 46] and absorption, Fig. 1 (f). Whisker-based Bi-2212 mesas have a turnstile antenna geometry [35], Fig. 2(b), enabling high radiation efficiency ($12\%$ at $f\simeq 4$ THz) [22] and causing the four-fold receiving directivity, Fig. 2 (f) [50]. Thus, the analysis of directivity (geometry-specific antenna effects) provides a valuable tool for investigating impedance matching and the efficiency of both absorption and emission.
@@
In conclusion, we have described the operational principle of a cascade switching current detector, utilizing arrays of coupled Josephson junctions. Coupling between junctions leads to the current locking phenomenon and results in avalanche-like switching of many junctions. This amplifies the readout voltage and enhances detector sensitivity. We have demonstrated two prototypes based either on low-$T_{c}$ linear Nb arrays, or high-$T_{c}$ stacked intrinsic Josephson junctions. Good impedance matching and absorption efficiency of the studied devices is confirmed by observation of geometry-specific antenna effects in receiving directivities. We conclude that Bi-2212 whisker-based cascade SCD are promising candidates for broad-band and high-performance THz sensors.
@@ Methods
The absorbed power, $P_{a}$, can be estimated from the suppression of the switching current, $\Delta I_{s}$, caused by the generated high-frequency current. In the MW case (low frequency limit) $\Delta I_{s}$ is just equal to the induced MW current [8]. $P_{a}$ is calculated using Eq. (4) from Ref. [19]: $P_{a}\simeq(2\sqrt{2}/3\pi)(\Delta I_{s}/I_{c0})^{3/2}I_{c0}V_{c}$. For the case of Fig. 2 (d), $\Delta I_{s}\simeq 3~{}\mu$A at the attenuation 3.7 dB, $I_{c0}=38~{}\mu$A, $V_{c}=20~{}$mV, yielding $P_{a}\simeq 5$ nW, approximately 30% of $P_{MW}$. The maximum voltage is $V\simeq 0.37$ V, and the sensitivity with respect to the absorbed power, $S_{a}=V/P_{a}\simeq 7.4~{}10^{7}$ (V/W) and the net sensitivity $S_{MW}=V/P_{MW}\simeq 2.6~{}10^{7}$ (V/W).
@@
The noise-equivalent power is determined as $NEP=v_{n}/S$, where $v_{n}$ is the voltage noise. It is typically in the range of few nV$/\sqrt{Hz}$ at room temperature, but for lock-in measurements at low $T$ it is certainly below 1 nV. Taking $v_{n}=1~{}$nV$/\sqrt{Hz}$ we obtain the following values of NEP with respect to absorbed power: $NEP_{MW}\simeq 1.35~{}10^{-17}$ W$/\sqrt{Hz}$ and $NEP_{THz}\simeq 6.25~{}10^{-17}$ W$/\sqrt{Hz}$.
@@ Acknowledgements
The work was supported by the Science for Peace and Security Programme, grant G5796. We are grateful to R. Gerdau and M. Galin for assistance with the fabrication of Nb-arrays and to A. Kalenyuk and E. Borodianskyi for assistance with the fabrication of Bi-2212 samples.